# Measuring the Temperature of Extremely Hot Shock-Heated Gas in the Major Merger MACS J0717.5++3745 With Relativistic Corrections to the Sunyaev-Zel’dovich Effect

### v1

@@ 5 Discussion
Combining the results of all three analyses, we find consistent average temperatures in the range of 15–20 keV for MACS J0717.5++3745. We firmly and independently establish that this major merger contains a significant amount of shock-heated gas from our rSZe analysis. In addition, other than the known calibration difference between XMM-Newton and Chandra, we find no evidence for differences in average temperature e.g. due to the specific distribution of gas density and temperature in MACS J0717.5++3745 or to potential biases from the band limitations of the X-ray facilities. While it may be tempting to utilize these measurements to better establish the temperature calibration of the X-ray satellites, we assert that such an inference is not justified due to the astrophysical and instrumental effects noted above, which are particularly relevant for an extremely hot merging system like MACS J0717.5++3745 even if they are not found to be statistically significant in our specific analysis. Further caution against such a comparison is warranted due to the lack of a homogeneous analysis of the Chandra and XMM-Newton data. Finally, we obtain a non-zero intrinsic scatter at modest statistical significance and with a characteristic value of 2–5 keV from all three probes. This suggests that the ICM in MACS J0717.5++3745 is not isothermal, and the variations in temperature over the projected face of the cluster are likely attributable to the ongoing multiple merger.
@@
While we caution against utilizing this present analysis as a tool to calibrate X-ray temperature measurements, we note that the rSZe does hold significant promise for obtaining such a calibration assessment. For instance, Biffi et al. (2014) find an on-average negligible bias between ICM temperatures obtained from different probes when restricting to morphologically regular clusters. Thus, an analysis based on a large sample of regular clusters would largely eliminate the potential astrophysical systematics related to ICM sub-structures noted above. In addition, the temperature of the ICM in such clusters should fall within the energy bands of the X-ray satellites.
@@
Looking to future spectrophotometric measurements of the SZe, this work has yielded some interesting insights. First, while FTS systems offer high spectral resolving power and potentially large mapping speeds, they are also very sensitive to mechanical and thermal changes that impact the optical path length while acquiring data. The SPIRE-FTS in particular was not designed to be sensitive to slowly varying continuum signals, and many of the low-frequency systematic effects in the data appear to be at least in part due to changes in the thermal properties of the instrument or the characteristics of the mechanical scanning mechanism. Optical systematic errors are also problematic in these measurements, with sensitivity to the precise temperature and configuration of the cold optics resulting in significant data loss. The inclusion of null data in the Dark Sky field to characterize these effects was crucial to extracting a scientific result from these observations. Secondly, as is made clear in Fig. 5, measurements of the rSZe should be optimized for bandwidths of $\sim 50\,$GHz ($R\sim 15$) at sub-mm wavelengths. This would optimize sensitivity to the broad continuum signal produced by the SZe, while allowing sensitivity to the spectral features in the rSZe deviation that provide contrast to statistically constrain the temperature of the ICM. Finally, this case study highlights the need for accurate information about the sub-mm galaxies in the cluster field. The rSZe effect would have been buried beneath the CIB sources if we did not have high-sensitivity broadband information to subtract their contribution to the surface brightness.
While this very deep space-based FTS measurement is likely to remain unique for the foreseeable future, we did successfully extract the faint SZe signal from it. This demonstrates that interesting new measurements of the ICM using the rSZe effect are well within range of a sensitive next-generation instrument with carefully controlled systematics and modest spectral capability.
@@ Acknowledgements
The scientific results reported in this article are based in part on observations made by the Chandra X-ray Observatory.
Herschel, CSO, XMM, Chandra, Keck

### v2

@@ V.2 Comparing rSZe and X-ray Measurements
Combining the results of all three analyses, we find consistent average temperatures in the range of 15–20 keV for MACS J0717.5++3745. We firmly and independently establish that this major merger contains a significant amount of shock-heated gas from our rSZe analysis. In addition, other than the known calibration difference between XMM-Newton and Chandra, we find no evidence for differences in average temperature e.g., due to the specific distribution of gas density and temperature in MACS J0717.5++3745 or to potential biases from the band limitations of the X-ray facilities. While it may be tempting to utilize these measurements to better establish the temperature calibration of the X-ray satellites, we assert that such an inference is not justified due to the astrophysical and instrumental effects noted above, which are particularly relevant for an extremely hot merging system like MACS J0717.5++3745 even if they are not found to be statistically significant in our specific analysis. Further caution against such a comparison is warranted due to the lack of a homogeneous analysis of the Chandra and XMM-Newton data. Finally, we obtain a nonzero intrinsic scatter at modest statistical significance and with a characteristic value of 2–5 keV from all three probes. This suggests that the ICM in MACS J0717.5++3745 is not isothermal, and the variations in temperature over the projected face of the cluster are likely attributable to the ongoing multiple merger.
@@
While we caution against utilizing this present analysis as a tool to calibrate X-ray temperature measurements, we note that the rSZe does hold significant promise for obtaining such a calibration assessment. For instance, Biffi et al. (2014) find an on-average negligible bias between ICM temperatures obtained from different probes when restricting to morphologically regular clusters. Thus, an analysis based on a large sample of regular clusters would largely eliminate the potential astrophysical systematics related to ICM substructures noted above. In addition, the temperature of the ICM in such clusters should fall within the energy bands of the X-ray satellites.
V.3 Prospects for Future Spectroscopic Measurements of rSZe
@@
Looking to future spectrophotometric measurements of the SZe, this work has yielded some interesting insights. First, while FTS systems offer high spectral resolving power and potentially large mapping speeds, they are also very sensitive to mechanical and thermal changes that impact the optical path length while acquiring data. The SPIRE-FTS, in particular, was not designed to be sensitive to slowly varying continuum signals, and many of the low-frequency systematic effects in the data appear to be at least in part due to changes in the thermal properties of the instrument or the characteristics of the mechanical scanning mechanism. Optical systematic errors are also problematic in these measurements, with sensitivity to the precise temperature and configuration of the cold optics resulting in significant data loss. The inclusion of null data in the Dark Sky field to characterize these effects was crucial to extracting a scientific result from these observations. Secondly, as is made clear in Fig. 5, measurements of the rSZe should be optimized for bandwidths of $\sim 50\,$GHz ($R\sim 15$) at submillimeter wavelengths. This would optimize sensitivity to the broad continuum signal produced by the SZe, while allowing sensitivity to the spectral features in the rSZe deviation that provide contrast to statistically constrain the temperature of the ICM. Finally, this case study highlights the need for accurate information about the submillimeter galaxies in the cluster field. The rSZe effect would have been buried beneath the CIB sources if we did not have high-sensitivity broadband information to subtract their contribution to the surface brightness.
V.4 Conclusion
We were able to measure the rSZe signal from MACS J0717.5++3745 at seven unique locations with the SPIRE Fourier transform spectrometer. These measurements were then aggregated by fitting a characteristic mean and intrinsic scatter, where we find an rSZe inferred average temperature of $T_{\rm rSZe}$ $15.1\begin{subarray}{c}+3.8\\ -3.3\end{subarray}$ keV which is consistent with X-ray inferred temperatures from both Chandra and XMM-Newton. While this very deep space-based FTS measurement is likely to remain unique for the foreseeable future, we have demonstrated that interesting new measurements of the ICM using the rSZe effect are well within the range of a sensitive next-generation instrument that has carefully controlled systematics and modest spectral capability. Measurements of this nature offer a complementary probe of both high-temperature and high-redshift ICM gas, which would allow for detailed study of the evolution of cluster thermodynamics that would otherwise be difficult or nonpermissible to understand utilizing other methods.
@@ Acknowledgements
This paper employs a dataset, obtained by the Chandra X-ray Observatory, contained in the Chandra Data Collection  https://doi.org/10.25574/cdc.548 (catalog DOI: 10.25574/1655).
Herschel, CSO, XMM, CXO, Keck I, and Keck II